# On action rate admissibility criteria

### v4

@@ 1 Introduction
The paper [9] introduces a version of the least action principle as a selection criterion in initial value problems possessing non-unique solutions. Interest arises from examples that use convex integration techniques to obtain families of non-unique wild solutions to the Euler system (De Lellis and Székelyhidi [4, 5]), and non-unique solutions to the Navier-Stokes equations (Buckmaster and Vicol [1]). Convex integration is also employed by Chiodaroli and Kreml [2] to investigate the two-dimensional Riemann initial value problem for the isentropic compressible Euler system. They construct a large set of entropic wild solutions to the Riemann problem. Consequently, in these problems it is desirable to look for an admissibility criterion that distinguishes a particular solution from the many possible. Because of its mathematical simplicity, we are motivated to seek an admissibitlity criterion that selects the classical 2-shock solution.
@@ 2 Least action admissibility criteria
where $F({\boldsymbol{u}})$ is a given operator and $t_{0},\,T]$ is the maximal time interval of existence of solutions belonging to $S$. Global existence is not assumed so that possibly $T<\infty.$. Later, the system (2.1) is specialised to be the Euler-Lagrange equation derived from a least action principle.
@@ 3.1 Convex integration
By corrugation we mean an iterative scheme of superposing on the sub-solution a sequence of spatio-temporal oscillations at decreasingly small scales such that in the limit the combined subsolution and oscillation converges to an appropriately defined weak (exact) solution to the equation of interest. In Nash’s fundamental paper [15] on isometric embedding of Riemannian manifolds, sub-solutions correspond to short embeddings. Nash produced a corrugation algorithm which he termed stages. Gromov [11, 12] generalised the technique, while De Lellis and Székelyhidi [4, 5] applied the general theory of convex integration to construct wild solutions to the Euler equations. The lecture notes by Markfelder [13] may be consulted for details. An overview is provided in the survey by De Lellis and Székelyhidi [6].
@@ 3.3 Fan sub-solutions
The next task is to define the sub-solution of Chiodaroli and Kreml [2, Sect. 3]. Together with the corrugations of Lemma 3.1 below it is an essential component of obtaining weak solutions by the convex integration process.
@@ 3.4 Corrugations
In their article Chiodaroli and Kreml [2] prove non-uniqueness of weak solutions to the Riemann initial value problem with initial conditions (3.5), i.e. not only is there the classical 2-shock solution given in Section 3.2, but also an infinite number of weak solutions that possess non-trivial $x_{1}$ dependence. Their proof, based upon convex integration, consists of two pieces: (i) a base sub-solution and (ii) an infinite sequence of corrugations which when combined with (i) yields an infinite number of solutions to the initial value problem. Here we recall these two concepts within the context of our Riemann initial value problem for the isentropic Euler equations. We emphasize that the corrugations combined with even a single sub-solution yield an infinite number of weak solutions to initial value problem with initial conditions (3.5).
@@ 5 Application of LAAP0
We conclude that $D(\rho_{m},t)<0$ for $t>0$. By continuity of $D(\rho_{1},t)$ we have $D(\rho_{1},t)<0$ for $t>0$ when $\rho_{1}$ belongs to a sufficiently small left neighborhood of $\rho_{m}$. Further, because at $t=0$ the domain of integration shown in Figure 1, where the integrand is non-zero, collapses to a set of measure zero, we have
@@
Given Riemann data satisfying (3.6), let $\varepsilon_{2}>0$ and assume that $\rho_{1}$ lies in a sufficiently small left neighborhood $I$ of $\rho_{m}$. Then the 2-shock solution is sARAC-admissible, and therefore is the strictly LAAP0-admissible solution in $S$.
@@
As just shown, for $\rho_{1}$ in the interval $I$ the action and its first and second time derivatives are lower for the 2-shock solution than for any other solution in $S$, including those obtained in [14] since for small enough time these coincide with those of [2]. Hence, both sARAC and strict LAAP0 exclude the admissibility of the convex integration and hybrid solutions and establish the strict admissibility of the 2-shock solution in $S$. ∎

### v5

@@ 1 Introduction
The paper [9] introduces a version of the least action principle as a selection criterion in initial value problems possessing non-unique solutions. Interest arises from examples that use convex integration techniques to obtain families of non-unique wild solutions to the Euler system (De Lellis and Székelyhidi [4, 5]), and non-unique solutions to the Navier-Stokes equations (Buckmaster and Vicol [1]). Convex integration is also employed by Chiodaroli and Kreml [2] to investigate the two-dimensional Riemann initial value problem for the isentropic compressible Euler system. They construct a large set of entropic wild solutions to the Riemann problem. Consequently, in these problems it is desirable to look for an admissibility criterion that distinguishes a particular solution from the many possible. Because of its mathematical simplicity, we are motivated to seek an admissibility criterion that selects the classical 2-shock solution.
@@ 2 Least action admissibility criteria
where $F({\boldsymbol{u}})$ is a given operator and $t_{0},\,T]$ is the maximal time interval of existence of solutions belonging to $S$. Global existence is not assumed so that possibly $T<\infty$. Later, the system (2.1) is specialised to be the Euler-Lagrange equation derived from a least action principle.
@@ 3.1 Convex integration
By corrugation we mean an iterative scheme of superposing on the sub-solution a sequence of spatio-temporal oscillations at decreasingly small scales such that in the limit the combined subsolution and oscillation converge to an appropriately defined weak (exact) solution to the equation of interest. In Nash’s fundamental paper [15] on isometric embedding of Riemannian manifolds, sub-solutions correspond to short embeddings. Nash produced a corrugation algorithm which he termed stages. Gromov [11, 12] generalised the technique, while De Lellis and Székelyhidi [4, 5] applied the general theory of convex integration to construct wild solutions to the Euler equations. The lecture notes by Markfelder [13] may be consulted for details. An overview is provided in the survey by De Lellis and Székelyhidi [6].
@@ 3.3 Fan sub-solutions
The next task is to define the sub-solution of Chiodaroli and Kreml [2, Sect. 3]. Together with the corrugations of Lemma 3.1 it is an essential component of obtaining weak solutions by the convex integration process.
@@ 3.4 Corrugations
In their article Chiodaroli and Kreml [2] prove non-uniqueness of weak solutions to the Riemann initial value problem with initial conditions (3.5), i.e. not only is there the classical 2-shock solution given in Section 3.2, but also an infinite number of weak solutions that possess non-trivial $x_{1}$ dependence. Their proof, based upon convex integration, consists of two pieces: (i) a base sub-solution and (ii) an infinite sequence of corrugations which when combined with (i) yields an infinite number of solutions to the initial value problem. Here we recall these two concepts within the context of our Riemann initial value problem for the isentropic Euler equations. We emphasise that the corrugations combined with even a single sub-solution yield an infinite number of weak solutions to initial value problem with initial conditions (3.5).
@@ 5 Application of LAAP0
We conclude that $D(\rho_{m},t)<0$ for $t>0$. By continuity of $D(\rho_{1},t)$ we have $D(\rho_{1},t)<0$ for $t>0$ when $\rho_{1}$ belongs to a sufficiently small left neighbourhood of $\rho_{m}$. Further, because at $t=0$ the domain of integration shown in Figure 1, where the integrand is non-zero, collapses to a set of measure zero, we have
@@
Given Riemann data satisfying (3.6), let $\varepsilon_{2}>0$ and assume that $\rho_{1}$ lies in a sufficiently small left neighbourhood $I$ of $\rho_{m}$. Then the 2-shock solution is sARAC-admissible, and therefore is the strictly LAAP0-admissible solution in $S$.
@@
As just shown, for $\rho_{1}$ in the interval $I$ the action and its first time derivative are less than or equal and the second time derivative strictly less for the 2-shock solution than for any other solution in $S$, including those obtained in [14] since for small enough time these coincide with those of [2]. Hence, both sARAC and strict LAAP0 exclude the admissibility of the convex integration and hybrid solutions and establish the strict admissibility of the 2-shock solution in $S$. ∎